OpenStudy (anonymous):

solve the equation 6^x+4(6 ^-x)=5

OpenStudy (loser66):

\[6^x + \frac{4}{6^x}=5\\let~6^x=u~I~have~u +\frac{4}{u}=5\\solve~for~u~and~then~plug~it~back~to~solve~for~x~by~log~property\]